1. Michaels Company’s total material costs are $50,000 and total conversion costs are $65,000. Equivalent units of production for materials are 10,000, and 5,000 for conversion costs. Instructions

Compute the unit costs for materials, conversion costs, and total manufacturing costs for the month. Show your answer in good form. Include all your calculations.

2.  

Sonys Inc. leases access to high-speed computers to small businesses. It provides the following information for the year:

      Budgeted           Actual

            Overhead cost                        $2,000,000      $1,900,000

            Computer hours                           100,000             90,000

            Direct labor hours                        200,000           180,000

Overhead is applied on the basis of computer hours.

Instructions

a. Compute the predetermined overhead rate.

b. Determine the amount of overhead applied for the year.

Show your answers in good form. Include all your calculations.
Answer #1

Unit cost = Total cost/Equivalent number of Units

Material = 50,000/10,000 = $5 per unit

Conversion cost = 65,000/5,000 = $13 per unit

Total Manufacturing cost = 5+13 = $18 per unit

a.Predetermined overhead rate = Budgeted Overheads/Budgeted Base

= 2,000,000/100,000

= $20 per computer hour

b.Overhead Applied = Predetermined Rate*Actual Hours

= 20*90,000

= $1,800,000